Solve:
12t-36=t^2
Please show me step by step to get the answer. Thanks

Answer by stanbon(75887)

12t-36=t^2
---
t^2 - 12t + 36 = 0
---
= (t-6)(t-6) = 0
---
= t = 6 with multiplicity 2
==============================
Cheers,
Stan H.
===========